Each of the equal sides of an isosceles triangle is 5cm greater that the base. The perimeter is 46cm. What is the lengh of each of the sides of the triangle?
Answer by checkley71(8403)

S+S+B=46 & S=B+5
B+5+B+5+B=46
3B=46-10
3B=36
B=36/3
B=12 CM.
THEN THE SIDES MUST BE
12+5=17
PROOF
17+17+12=46
46=46